OpenStudy (anonymous):

What is the slope of the line passing through the points (-3,4) and (2,-1)? A) 1 B) -1 C) 3/5 D) -5/3

OpenStudy (anonymous):

@Monkey4224

OpenStudy (anonymous):

i think its d?

OpenStudy (anonymous):

and it doesnt come with a picture (just to let you know)

OpenStudy (anonymous):

no idea.@pooja195 @triciaal @Michele_Laino @Kainui

OpenStudy (freckles):

evaluate \[\frac{y_1-y_2}{x_1-x_2} \\ \text{ where } (x_1,y_1)=(-3,4) \text{ and } (x_2,y_2)=(2,-1)\]

OpenStudy (anonymous):

ohh ok let me try the work

OpenStudy (anonymous):

what is 2-(-3)?

OpenStudy (freckles):

2-(-3) is the same as 2+3 since - times a - is positive

OpenStudy (anonymous):

ok so that means that i got -1! thx

OpenStudy (anonymous):

@freckles btw your expression i had to do it backwards so that i made sense

OpenStudy (freckles):

\[\frac{4-(-1)}{-3-2} \text{ is the same as } \frac{-1-4}{2-(-3)} \\ \] either way gives you the correct slope

OpenStudy (freckles):

either way makes sense that is \[\frac{y_2-y_1}{x_2-x_1} \text{ is the same as computing } \frac{y_1-y_2}{x_1-x_2}\]

OpenStudy (anonymous):

oh cuz my math teacher told me to do it backwards from the original way you wrote it but any way thx

OpenStudy (anonymous):

i fan btw and medal too